Question 1.

The management of Zenith plc are to assemble a new computer aimed at the business market. The forecast cost structure for producing this new computer is as follows:

Materials Component A

$150 per unit.

Materials Component B

$300 per unit.

Labour

Labour is paid on a piecework basis of $100 per unit.

Fixed Overheads

For activity between zero and 80,000 units per annum the expected cost is $5.8 million.

Sales

It is expected that the selling price will be $710 per computer.

Required:

  1. Determine the profit when producing 0, 30,000, 40,000, and 80,000 units per annum.

  2. Calculate the break-even point.
Answer #1
Output 0 Units Output 30,000 Units Output 40,000 Units Output 80,000 Units
Sales(1) 0 30,000 x 710 =21,300,000 40,000 x 710 =28,400,000 80,000 x 710 =56,800,000
Expenses:
Material 0 30,000 x 450 = 13,500,000 40,000 x 450 = 18,000,000 80,000 x 450 = 36,000,000
Labor 0 30,000 x 100 = 3,000,000 40,000 x 100 = 4,000,000 80,000 x 100 = 8,000,000
Fixed overhead 5,800,000 5,800,000 5,800,000 5,800,000
Total Expenses (2) 5,800,000 22,300,000 27,800,000 49,800,000
Profit /Loss(1-2) -5,800,000 -1,000,000 6,000,000 7,000,000

Material per unit = Cost of materials A + Cost of material B

= 150+300

= $450

Variable cost per unit = Direct materials + Direct labor

= 450+100

= $550

Selling price per unit = $710

Contribution margin per unit = Selling price per unit - Variable cost per unit

= 710-550

= $160

Break even point = Fixed cost/ Contribution margin per unit

= 5,800,000/160

= 36,250

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks